OpenStudy (anonymous):

After picking apples, 4 friends compared the amount of apples in their buckets. Dan’s bucket was 2/3 full, Mary’s bucket was 4/5 full, Julie’s bucket was 3/4 full, and Jane’s bucket was 1/8 full. How would these amounts be arranged in order from greatest to least?

OpenStudy (anonymous):

In order to order these from greatest to least, you must have a common denominator. You have your fractions: 2/3, 4/5, 3/4, 1/8. Now you have to find a common multiple of 3, 5, 4, and 8. How about 120? I know it is large, but it is the smallest common multiple of the four numbers. In order to get 2/3 to ?/120, you must multiply the numerator and denominator by 40. 2*40=80 3*40=120. So your fraction is 80/120. 4/5 to ?/120, multiply both by 24. 4*24=96 5*24=120. So your fraction is 96/120. 3/4 to ?/120 multiply both by 30. 3*30=90 4*30=120. So your fraction is 90/120. 1/8 to ?/120 multiply both by 15. 1*15=15 8*15=120. So your fraction is 15/120. Now you have all of your fractions:80/120(2/3), 96/120(4/5), 90/120(3/4), 15/120(1/8). Since they all have the same denominator, we just look at the numerators. Which number is the greatest out of 80, 96, 90, and 15? 96. Then you continue to order them from greatest to least.... 96, 90, 80, 15. So your fractions would be.. 96 90 80 15 -- -- -- -- 120 120 120 120 Looking back, which fraction matched each fraction? 4 3 2 1 - - - - 5 4 3 8 Now that is your order :) Hope it helped explain what you needed!